Write an equation of the line containing the given point and parallel to the given line. Express the answer in the form y=mx+b (7,9); 7x+y=2. The equation of the line is y=. I have tried solving this and each time I get a different answer. Now I'm confused.
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Write an equation of the line containing the given point and parallel to the given line. Express the answer in the form y=mx+b (7,9); 7x+y=2.
----------
Note: The answer is not y=1
===============================
Find the slope of the given line:
y = -7x + 2
slope = -7
-------------------
Any line parallel to that given line must have a slope of -7
The line you want has a slope of -7 and passes thru y=9 when x=7:
Solve for "b":
9 = -7*7 + b
9 = -49 + b
b = 58
==============
Equation:
y = -7x + 58
=====================
Cheers,
Stan H.

Answer by richwmiller(17219)

So have you looked for similar problems or lessons ?
No of course not.
What have you done? You said you tried several attempts but without telling us anything.